Kofinatland is a hypothetical economy with the following
economic indicators:
. Size of population: 8 million
. GDP at market price (GDP at m.px.): US $570 billion
. Subsidies: US $50 billion

. Interest paid by consumers: US $27 billion
Employees contribution to social insurance: US $200
billion
Net factor income from abroad (NFYFA): US $200
billion

Capital consumption allowance (CCA): US $80 billion
. Human development index (HDI): 0.91
Indirect taxes on goods and services (ITGS): US $120
billion
. Life expectancy rate: 87 years

. Government transfer payment: US $71.5 billion
Factor income of foreigners in the economy (FYF): US
$80 billion
. Natural population growth rate: 0.5%.

Determine the economy's:
i. Net national product at market cost (NNP at m.px)
ii. Net national product at factor cost (NNP at f.c.)
ii. Factor income of nationals abroad (FYN).
iiii. GDP per capita (note to use the GDP at m. px. in your
1
Expert's answer
2020-05-08T14:23:58-0400

i.Net national product at market cost=GNP-Capital consumption allowance=GDP at market price (GDP at m.px.)-Net factor income from abroad-Capital consumption allowance=570-200-80=290

ii.Net national product at factor cost (NNP at f.c.) =GNP at market price -Indirect taxes on goods and services-Capital consumption allowance=570-120-80=370

iii.Factor income of nationals abroad (FYN)=NDP at MP + Depreciation + Factor income from abroad - GNP at FC - Indirect Taxes + Subsidies=570-80+80+200-570-120+50=130

iiii.GDP per capita (note to use the GDP at m. px.)=GDP at market price (GDP at m.px.)/Size of population:=570/8=71.25